OpenStudy (anonymous):

i need some reviewing on elimination. -4x - 2y = -12 4x + 8y = -24 prefer if steps can be shown and possibly explain.

OpenStudy (amistre64):

elimination is when we add the two equations together and one of the variables goes away

OpenStudy (amistre64):

if we add these 2 together, do you see a variable that goes away?

OpenStudy (angela210793):

add both sides--> 6y=-36 y=-6 replace y=-6 in one of the eq--> 4x=-24+48 4x=24 x=6 solution (6;-6)

OpenStudy (anonymous):

4x goes away

OpenStudy (amistre64):

that right, the x part goes away leaving us with only the y part to solve for :)

OpenStudy (anonymous):

well, i am a little stuck on the substituting part. why wouldn't it be 4x = -24 -48. because -6*8 is negative

OpenStudy (anonymous):

add the two equations in a column. \[-4x+4x=0\] \[-2y+8y=6y\] \[-12+-24=-36\]

OpenStudy (anonymous):

now you have \[6y=-36\] so \[y=\frac{-36}{6}\] \[y=-6\]

OpenStudy (anonymous):

now y is not a mystery or a variable, it is known to be -6, so the first equation is \[-4x-2(-6)=-12\] and so \[-4x+12=-12\] subtract 12 from both sides to get \[-4x=-24\] divide by -4 to get \[x=\frac{-24}{-4}=6\]

OpenStudy (amistre64):

elimination is another way to solve for the system of equations; "substitution" in general is not needed as a result

OpenStudy (anonymous):

but i see this is solved earlier, so not sure what i added if anything

OpenStudy (amistre64):

you added blood sweat and tears :)

OpenStudy (anonymous):

you are showing your age!

OpenStudy (anonymous):

thanks for helping. i got it now. you cleared things up.